# Coloring Groups

### v1

@@ 1 Introduction
Cameron and Fon-der-Flaass first introduced toggle groups on order ideals [2] in order to study the map now known as rowmotion (the name “toggle group” was introduced later in [10]). Given a poset $P$, each element $p\in P$ has a corresponding permutation $\tau_{p}$ (called a toggle) of the order ideals $\mathcal{J}(P)$ of $P$ which acts in the following way, toggling the element $p$ as long as the result is still an order ideal.
@@
Definition 1.1 (Toggle Group).
[9]
@@
The idea of order ideal toggling has been generalized from the purely combinatorial perspective to the piece-wise linear and birational realms in [3] and  [5, 6] respectively. These groups were studied in-depth by Striker in [9] when arising from some families of combinatorial and geometric objects. More recently in [1] they have been considered with respect to their group theoretic properties such as primitivity.
@@
Acknowledgements: The authors would like to thank Jonathan Bloom and Jessica Striker for their guidance and many, many helpful conversations. They would also like to thank Nate Lesnevich for his insights on an early version of this project.
@@ 2 Preliminaries and Definitions
$\mathfrak{G}_{\kappa}=\langle{\color[rgb]{0,0.5234375,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5234375,1}\pgfsys@color@rgb@stroke{0}{0.5234375}{1}\pgfsys@color@rgb@fill{0}{0.5234375}{1}(2,3)(4,5)},{\color[rgb]{1,0.56640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.56640625,0}\pgfsys@color@rgb@stroke{1}{0.56640625}{0}\pgfsys@color@rgb@fill{1}{0.56640625}{0}(3,4)(6,7)},{\color[rgb]{0.56640625,0.828125,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.56640625,0.828125,0}\pgfsys@color@rgb@stroke{0.56640625}{0.828125}{0}\pgfsys@color@rgb@fill{0.56640625}{0.828125}{0}(1,2)(3,6)}\rangle\cong GL(3,2).$
@@ 3 Tools for Analyzing Coloring Groups
    The final case is when $q=2$ and $d\geq 3$. If Inequality 4 does not hold, it is straightforward to show that $d\leq 8$. The only values between $3$ and $8$ for which $12^{d}-$ 1 is prime are $d=3,5,$ or $7$. When $d=3$, we have that $p=7$, so $\mathfrak{G}_{\kappa}$ must contain $S_{7}$ (See Appendix B). For $d=5,7$ we have that $|PGL_{d}(q)|>|PGL_{2}(p)|$.
@@
Let $G$ be a graph with proper edge coloring $\kappa:E(G)\to[k]$. Let $e$ be an edge such that every edge incident to $e$ is of a distinct color. We call $e$ a symmetric edge if there exists a subset $S$ of $[k]$ containing the colors of all edges incident to $e$ but not containing the color of $e$ such that removing all edges colored by a color in $S$ leaves a graph with exactly one even component.
@@ 4.1 Generalized Toggle Groups
Let $(G,\kappa)$ be a proper edge coloring of a tree. A necessary condition for $\mathfrak{G}_{\kappa}$ to be a generalized toggle group is that for every path, reading the edge colors along that path must be a toggle word.
@@
This follows from the following observation on toggle graphs. Let $G=P_{L}$ for some toggle group $T(L)$ and suppose that $w_{1}w_{2}\dots w_{\ell}$ is the sequence elements of the base set corresponding to the coloring for a walk in $G$ with no consecutive letters. Without loss of generality we can assume no vertices of $G$ are repeated.
@@
If exactly one label appears an odd number of times, call it $e$, then the result of this walk is that $e$ is either toggled in or out of the initial set, as all other labels are toggled in and out an even number of times. As no vertices are repeated then we could also have taken the distinct path of applying the toggle $\tau_{e}$, so $G$ must contain a cycle. This edge must exist as $G$ arose from a toggle group.
@@
The following edge coloring gives rise to a coloring group isomorphic to a generalized toggle group. The family of subsets are the vertex labels.
@@ 4.2 Independence Posets
When Cameron and Fon-der-Flaass originally described the toggle group, they were essentially considering it as a group structure on a distributive lattice. While the generalized toggle group is an extension of these order ideal toggles from a group-theoretic perspective to subsets of a finite set, there has historically been considerably more focus on trying to understand generalizations of this family of lattices from a combinatorial as well as order-theoretic point of view. One particular family that serves as a generalization are the trim lattices, first introduced by Hugh Thomas in [11]. Before we define these lattices, recall that an element $x$ of a lattice $L$ is left modular if for any $y<z\in L$ the following equality holds $y\vee(x\wedge z)=(y\vee x)\wedge z$.
Definition 4.6.
[11] A lattice is trim if it has a maximal chain of $(n+1)$ left modular elements as well as $n$ join-irreducible elements and $n$ meet irreducible elements.
@@
as well as many more. To be self-contained, we now follow the treatment of  [12] to define these objects, up to the omission of proofs.
For the remainder of the background on independence posets, let $G$ be a directed acylic graph. A $G$-order is the partial order on $V(G)$ where $g_{1}\geq g_{2}$ when there is a directed path from $g_{1}$ to $g_{2}$. Recall that an independent set of $G$ is a subset where no two elements are adjacent. To define an independence poset we first need the following definitions.
Definition 4.7.
[12, Definition 1.2] A pair ($D$, $U$) of independent sets of $G$ is called orthogonal if there is no edge in $G$ from an element of $D$ to an element of $U$. An orthogonal pair of independent sets ($D$, $U$) is called tight if whenever any element of $D$ is increased (removed and replaced by a larger element with respect to the $G$-order), any element of $U$ is decreased, or a new element is added to either $D$ or $U$, then the result is no longer an orthogonal pair of independent sets. We abbreviate tight orthogonal pair by top, and we write $\textbf{top}(G)$ for the set of all tops of $G$.
@@
Theorem 4.8.
[12, Theorem 1.2] Let $I$ be an independent set of a directed acyclic graph $G$. Then there exists a unique $(I,U)\in\textbf{top}(G)$ and a unique $(D,I)\in\textbf{top}(G)$.
@@
Definition 4.9.
[12, Definition 1.3] The flip of $(D,U)\in\textbf{top}(G)$ at an element $g\in G$ is the tight orthogonal pair $flip_{g}(D,U)$ defined as follows: if $g\notin D$ and $g\notin U$, the flip does nothing. Otherwise, preserve all elements of $D$ that are not less than $g$ and all elements of $U$ that are not greater than $g$ (and delete all other elements); after switching the set to which $g$ belongs, then greedily complete $D$ and $U$ to a tight orthogonal pair in the orders $\ell^{\prime}$ and $\ell$, respectively.
@@
Definition 4.10.
[12, Definition 1.4] The independence poset on $\textbf{top}(G)$ is the reflexive and transitive closure of the relations $(D,U)\lessdot(D^{\prime},U^{\prime})$ if there is some $g\in U$ such that $flip_{g}(D,U)=(D^{\prime},U^{\prime}).$
@@
Definition 4.14.
[9, Definition 2.5] Given $L\subseteq 2^{E}$, let an essentialization of $(L,E)$ be $(L^{\prime},E^{\prime})$ where $E^{\prime}$ and $L^{\prime}$ are as follows. Every $e\in E^{\prime}$ must be an element $e\in E$ such that there exist $X,Y\in L$ with $e\in X$ and $e\not\in Y$. That is, no element of $E^{\prime}$ may be contained in all or none of the subsets in $L$. Also, consider subsets $Z\subseteq E$ with $|Z|>1$ such that if $e\in Z$ and $e\in W\in L$, then $Z\subseteq W$. For any maximal such $Z$, we exclude all elements of $Z$ except one from $E^{\prime}$. Let $L^{\prime}=\{X\cap E^{\prime}\ |\ X\in L\}$.
Definition 4.15.
[9, Definition 2.21] We define the collection of inductively toggle-alternating sets $L$ to be the smallest collection of $L\subseteq 2^{E}$, with essentializations $(L^{\prime},E^{\prime})$, such that:
@@
Theorem 4.16.
[9, Theorem 2.22] If $L$ is inductively toggle-alternating, then it is toggle-alternating.
@@
The proof is almost exactly the proof of 4.16 due to Striker in [9], which is essentially the proof in the case of order ideals due to Cameron and Fon-der-Flaass in  [2].
@@
Before proving Theorem 4.12, we mention a few more necessary results on the structure of independence posets from [12]. The first is the tight orthogonal pair recursion. For any $g\in G$, $\{g\}$ is an independent set, there are unique $D,U$ such that $(D,\{g\})=m_{g}$ and ($(\{g\},U)=j_{g}$ are tight orthogonal pairs. Let $\textbf{top}_{g}(G)=[\hat{0},m_{g}]$ and $\textbf{top}^{g}(G)=[j_{g},\hat{1}]$. In the case where $g$ is an extremal (minimal or maximal) element of the $G$-order, these two intervals in fact partition the independence poset. Formally this is stated as follows.
Lemma 4.19.
[12, Lemma 3.7] Let $G$ be an acyclic directed graph. If $g$ is an extremal element of $G$ then $\textbf{top}(G)=\textbf{top}_{g}(G)\cup\textbf{top}^{g}(G)$. Furthermore,
@@
Lemma 4.20.
[12, Lemma 3.9] Let $G$ be a directed acyclic graph.
@@
Theorem 4.21.
[12, Theorem 3.10] Let $g$ be an extremal element of an acyclic directed graph $G$. Then

### v4

@@ Abstract
keywords:
Groups from graphs, toggle group, independence poset
@@ 1 Introduction
Cameron and Fon-der-Flaass in  [2] first introduced toggle groups on order ideals in order to study the map now known as rowmotion (the name “toggle group” was introduced later in [9]). Given a poset $P$, each element $p\in P$ has a corresponding permutation $\tau_{p}$ (called a toggle) of the order ideals $\mathcal{J}(P)$ of $P$ which acts in the following way, toggling the element $p$ as long as the result is still an order ideal.
@@
Definition 1.1 (Toggle Group [8]).
@@
The idea of order ideal toggling has been generalized from the purely combinatorial perspective to the piece-wise linear and birational realms in [3] and [6, 5] respectively. These groups were studied in-depth by Striker in [8] when arising from some families of combinatorial and geometric objects. More recently, by Bloom and Saracino in [1], they have been considered with respect to their group theoretic properties such as primitivity.
@@ 2 Preliminaries and Definitions
$\mathfrak{G}_{\kappa}=\langle{\color[rgb]{0,0.5234375,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5234375,1}(2,3)(4,5)},{\color[rgb]{1,0.56640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.56640625,0}(3,4)(6,7)},{\color[rgb]{0.56640625,0.828125,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.56640625,0.828125,0}(1,2)(3,6)}\rangle\cong GL(3,2).$
@@ 3 Tools for Analyzing Coloring Groups
    The final case is when $q=2$ and $d\geq 3$. If Inequality 2 does not hold, it is straightforward to show that $d\leq 8$. The only values between $3$ and $8$ for which $12^{d}-$ 1 is prime are $d=3,5,$ or $7$. When $d=3$, we have that $p=7$, so $\mathfrak{G}_{\kappa}$ must contain $S_{7}$ (See Appendix B). For $d=5,7$ we have that $|PGL_{d}(q)|>|PGL_{2}(p)|$.
@@
Let $G$ be a graph with proper edge coloring $\kappa:E(G)\to[k]$. Let $e$ be an edge such that all edges incident to $e$ are of distinct colors. We call $e$ a symmetric edge if there exists a subset $S$ of $[k]$ containing the colors of all edges incident to $e$ but not containing the color of $e$ such that removing all edges colored by a color in $S$ leaves a graph with exactly one even component.
@@ 4.1 Generalized Toggle Groups
Let $(G,\kappa)$ be a proper edge coloring of a tree. A necessary condition for $\mathfrak{G}_{\kappa}$ to be a generalized toggle group is that for every path of length greater than one, reading the edge colors along that path must be a toggle word.
@@
This follows from the following observation on toggle graphs. Let $G=P_{L}$ for some toggle group $T(L)$ and suppose that $w_{1}w_{2}\dots w_{\ell}$ is the sequence of elements of the base set corresponding to the coloring for a walk in $G$ with no consecutive letters. Without loss of generality we can assume no vertices of $G$ are repeated.
@@
If exactly one label appears an odd number of times, call it $e$, then the result of this walk is that $e$ is either toggled in or out of the initial set $S$, as all other labels are toggled in and out an even number of times. That is, this path ends at $\tau_{e}(S)$. Because $G$ arose from a toggle group, there must be an edge labeled $e$ between $S$ and $\tau_{e}(S)$. This single edge constitutes a path distinct from $w_{1}w_{2}\dots w_{\ell}$ between $S$ and $\tau_{e}(S)$, so $G$ must contain a cycle.
@@
The following edge coloring gives rise to a coloring group isomorphic to a generalized toggle group. The subsets are the vertex labels.
@@ 4.2 Independence Posets
When Cameron and Fon-der-Flaass originally described the toggle group, they were essentially considering it as a group structure on a distributive lattice. While the generalized toggle group is an extension of these order ideal toggles from a group-theoretic perspective to subsets of a finite set, there has historically been considerably more focus on trying to understand generalizations of this family of lattices from a combinatorial as well as order-theoretic point of view. One particular family that serves as a generalization is that of trim lattices, first introduced by Thomas in [11]. Before we define these lattices, recall that an element $x$ of a lattice $L$ is left modular if for any $y<z\in L$ the following equality holds $y\vee(x\wedge z)=(y\vee x)\wedge z$.
Definition 4.6 ([11]).
A lattice is trim if it has a maximal chain of $(n+1)$ left modular elements as well as $n$ join-irreducible elements and $n$ meet irreducible elements.
@@
as well as many more. To be self-contained, we follow the treatment of  [12] to define these objects, up to the omission of proofs.
For the remainder of the background on independence posets, let $G$ be a directed acyclic graph. The associated $G$-order is the partial order on $V(G)$ where $g_{1}\geq g_{2}$ when there is a directed path from $g_{1}$ to $g_{2}$. Recall that an independent set of $G$ is a subset where no two elements are adjacent. To define an independence poset we first need the following definitions.
Definition 4.7 ([12, Definition 1.2]).
A pair ($D$, $U$) of independent sets of $G$ is called orthogonal if there is no edge in $G$ from an element of $D$ to an element of $U$. An orthogonal pair of independent sets ($D$, $U$) is called tight if whenever any element of $D$ is increased (removed and replaced by a larger element with respect to the $G$-order), any element of $U$ is decreased, or a new element is added to either $D$ or $U$, then the result is no longer an orthogonal pair of independent sets. We abbreviate tight orthogonal pair by top, and we write $\textbf{top}(G)$ for the set of all tops of $G$.
@@
Theorem 4.8 ([12, Theorem 1.2]).
Let $I$ be an independent set of a directed acyclic graph $G$. Then there exists a unique $(I,U)\in\textbf{top}(G)$ and a unique $(D,I)\in\textbf{top}(G)$.
@@
Definition 4.9 ([12, Definition 1.3]).
The flip of $(D,U)\in\textbf{top}(G)$ at an element $g\in G$ is the tight orthogonal pair $flip_{g}(D,U)$ defined as follows: if $g\notin D$ and $g\notin U$, the flip does nothing. Otherwise, preserve all elements of $D$ that are not less than $g$ and all elements of $U$ that are not greater than $g$ (and delete all other elements); after switching the set to which $g$ belongs, then greedily complete $D$ and $U$ to a tight orthogonal pair in the orders $\ell^{\prime}$ and $\ell$, respectively.
@@
Definition 4.10 ([12, Definition 1.4]).
The independence poset on $\textbf{top}(G)$ is the reflexive and transitive closure of the relations $(D,U)\lessdot(D^{\prime},U^{\prime})$ if there is some $g\in U$ such that $flip_{g}(D,U)=(D^{\prime},U^{\prime}).$
@@
Definition 4.14 ([8, Definition 2.5]).
Given $L\subseteq 2^{E}$, let an essentialization of $(L,E)$ be $(L^{\prime},E^{\prime})$ where $E^{\prime}$ and $L^{\prime}$ are as follows. Every $e\in E^{\prime}$ must be an element $e\in E$ such that there exist $X,Y\in L$ with $e\in X$ and $e\not\in Y$. That is, no element of $E^{\prime}$ may be contained in all or none of the subsets in $L$. Also, consider subsets $Z\subseteq E$ with $|Z|>1$ such that if $e\in Z$ and $e\in W\in L$, then $Z\subseteq W$. For any maximal such $Z$, we exclude all elements of $Z$ except one from $E^{\prime}$. Let $L^{\prime}=\{X\cap E^{\prime}\ |\ X\in L\}$.
Definition 4.15 ([8, Definition 2.21]).
Define the collection of inductively toggle-alternating sets $L$ to be the smallest collection of $L\subseteq 2^{E}$, with essentializations $(L^{\prime},E^{\prime})$, such that:
@@
Theorem 4.16 ([8, Theorem 2.22]).
If $L$ is inductively toggle-alternating, then it is toggle-alternating.
@@
The proof is almost exactly the proof of 4.16 due to Striker in [8], which is essentially the proof in the case of order ideals of Cameron and Fon-der-Flaass [2].
@@
Before proving Theorem 4.12, we mention a few more necessary results on the structure of independence posets from [12]. The first is the tight orthogonal pair recursion. For any $g\in G$, $\{g\}$ is an independent set, and there are unique $D,U$ such that $(D,\{g\})=m_{g}$ and ($(\{g\},U)=j_{g}$ are tight orthogonal pairs. Let $\textbf{top}_{g}(G)=[\hat{0},m_{g}]$ and $\textbf{top}^{g}(G)=[j_{g},\hat{1}]$. In the case where $g$ is an extremal (minimal or maximal) element of the $G$-order, these two intervals in fact partition the independence poset. Formally this is stated as follows.
Lemma 4.19 ([12, Lemma 3.7]).
Let $G$ be an acyclic directed graph. If $g$ is an extremal element of $G$ then $\textbf{top}(G)=\textbf{top}_{g}(G)\cup\textbf{top}^{g}(G)$. Furthermore,
@@
Lemma 4.20 ([12, Lemma 3.9]).
Let $G$ be a directed acyclic graph.
@@
Theorem 4.21 ([12, Theorem 3.10]).
Let $g$ be an extremal element of an acyclic directed graph $G$. Then
@@ 12 $(C_{3}\times C_{3})\rtimes((C_{2})^{4}\rtimes C_{2})$ 288 No 1,2,1,3,1,2,1,3,1,2,1
Acknowledgements.
The authors would like to thank Jonathan Bloom and Jessica Striker for their guidance and many, many helpful conversations. They would also like to thank Nate Lesnevich for his insights on an early version of this project.